a car travels at a speed of 40 km/h for 2 hours and then at 60 km/h for 3 hours.What is the average speed of the car during the entire journey?
Answer by lwsshak3(11628)

a car travels at a speed of 40 km/h for 2 hours and then at 60 km/h for 3 hours.What is the average speed of the car during the entire journey?
***
speed=distance/travel time
distance=2*40+60*3=80+180=260 km
travel time=2+3=5 hrs
Average speed during entire journey=260/5=52 km/h